OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8 A. one solution B. two solutions C. infinitely many solutions D. no solution

OpenStudy (anonymous):

A. one solution. there is one point that satisfies. if you want to check, graph both functions and find the intersect. that point, and that point only, will satisfy both equations.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

they are the same equation there are infinitely many solutions

OpenStudy (anonymous):

use substitution

OpenStudy (anonymous):

use nothing they are the same equation

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

x = -4y + 4 2x + 8y = 8 ----------------- 2x + 8y = 8 2(-4y + 4) = 8 -8y + 8 = 8 -8y = 0 --- --- -8 -8 y = 0 ---------------- x = -4y + 4 x = -4(0) + 4 x = 0 + 4 x = 5 ---------------------------- Solution (5, 0) 1 solution